# Aerodynamic and structural airfoil shape optimisation via Transfer Learning-enhanced Deep Reinforcement Learning

### v1

@@ Abstract
The main objective of this paper is to introduce a transfer learning-enhanced, multi-objective, deep reinforcement learning (DRL) methodology that is able to optimise the geometry of any airfoil based on concomitant aerodynamic and structural criteria. To showcase the method, we aim to maximise the lift-to-drag ratio while preserving the structural integrity of the airfoil –as modelled by its maximum thickness– and train the DRL agent using a list of different transfer learning (TL) strategies. The performance of the DRL agent is compared with Particle Swarm Optimisation (PSO), a traditional gradient-free optimisation method. Results indicate that DRL agents are able to perform multi-objective shape optimisation, that the DRL approach outperforms PSO in terms of computational efficiency and shape optimisation performance, and that the TL-enhanced DRL agent achieves performance comparable to the DRL one, while further saving substantial computational resources.
1 Introduction
@@
Airfoil design is key to improving the performance and efficiency of various aerodynamic systems, such as aircraft wings, turbine blades, and wind turbines [1]. Such design often predicates on (i) achieving an optimal balance between two aerodynamic forces –where one aims to maximise the aerodynamic lift and minimise the aerodynamic drag– while (ii) maintaining structural integrity [2]. The aerodynamic balance is often explicitly modelled as the lift-to-drag ratio , where and are the lift and drag coefficients, and the overall problem is mathematised in a so-called objective function , where denotes the set of design variables. Finding proper designs thus requires to find the values that optimise such objective function. Accordingly, airfoil design reduces to a shape optimisation problem.
@@
The two common strategies for airfoil shape optimisation involve, respectively, using gradient-based or gradient-free methods. The former approach utilise the gradient , to find the local directions that maximally increase the objective function and iteratively refines the parameters accordingly [3, 4]. This approach is computationally efficient –particularly with high-dimensional design spaces–, making it a popular choice in aerodynamics as highlighted by comparative studies [5, 6]. The well-known adjoint method, introduced in Fluid Dynamics by the seminal work of Pironneau [7], is one prominent example of a gradient-based technique; requiring a computational effort that scales independently of the number of design variables. This method was subsequently refined by the so-called discrete adjoint method, becoming a widely used choice in this field [8, 9]. While gradient-based methods are generally efficient and very popular in aerodynamic shape optimisation, they only have provable local convergence, and thus for non-convex objective functions their performance is sensitive to the initial design point [10] and function continuity [3].
@@
On the other hand, gradient-free methods offer an alternative to complex, nonlinear optimisation problems that often appear in aerodynamic design [3, 11]. Despite their higher computational cost compared to gradient-based methods, they are more robust at finding global optima and handling challenging objective functions [3, 4]. Popular gradient-free methods deployed in the context of airfoil shape optimisation
include Genetic Algorithms (GA) [12] or
@@
Particle Swarm Optimisation (PSO) [13, 14, 15]. As advanced, these methods are often limited by their computational demands and can be inefficient with very large design spaces [6]. To partially bridge these issues, some recent efforts to improve the computational efficiency of gradient-free methods include, e.g., parallel implementations, which can make GAs viable even for intensive applications such as those in aerospace engineering [16].
@@
Recently, deep reinforcement learning (DRL) [17] has emerged as a promising (gradient-free) tool for shape optimisation [18, 19]. DRL methods have indeed demonstrated remarkable capabilities in various domains, particularly when dealing with complex high-dimensional data [4]. Their impact in fluid dynamics studies and aeronautical applications indeed belongs to a recent yet rapidly expanding research field that encompasses both general-purpose ML-based fluid mechanics [20, 21, 22], and their application to several aeronautical problems [23, 24] with clear industrial impact [25].
@@
In the task of airfoil optimisation, a significant drawback of training DRL algorithms lies in the computational cost associated with solving the aerodynamic flow problem around the airfoil for a concrete shape: while the most accurate approach to solve the aerodynamic flow involves computational fluid dynamics (CFD) techniques, its intrinsic high computational cost requires opting for alternative, approximate solvers. For instance, some recent studies employ panel methods such as XFoil [18, 19] to approximate the aerodynamic performance of airfoils, while some other studies [26, 27] focus on further reducing optimisation costs by constructing surrogate models instead. Although the latter offers faster training times and smoother reward variations with respect to the agent’s actions –facilitating the learning process–, these approximate models come at the price of a decrease in accuracy.
@@
Interestingly, one significant advantage of recent deep learning methods in optimisation tasks is their potential to enhance their learning efficiency through so-called Transfer Learning [28, 29]: the possibility of pre-training a given model on one task and subsequently fine-tuning it for a related but different task. Transfer learning could also be developed in the context of DRL, accelerating the agent’s learning process by reusing elements such as policies or feature representations and effectively reducing the need for extensive exploration, a particularly valuable trait in aerodynamic optimisation.
To the best of our knowledge this approach has been seldom explored, with a few encouraging exceptions for missile control surface and airfoils [28, 30].
@@
Additionally, note that most studies on airfoil shape optimisation focus solely on maximising aerodynamic efficiency [26, 18, 19, 28, 31], without explicitly considering important structural constraints [32] related, e.g., to the airfoil’s area or thickness. Accordingly, there is a need to develop shape optimisation frameworks that aim to maximise the aerodynamic efficiency while preserving the structural integrity of the airfoil. This, we argue, can be interpreted as an example of multi-objective optimisation [26].
@@
To bridge these gaps, here we aim to explore the performance of DRL (enhanced with different possible Transfer Learning strategies) as a gradient-free framework for airfoil shape optimisation that considers the concomitant role of aerodynamic efficiency and structural integrity. To that aim, we define a flexible reward function that combines the lift-to-drag ratio of the flow around the airfoil and the airfoil’s maximum thickness. We initially train our DRL agent using XFoil –both for aerodynamic and combined aerodynamic/structural shape optimisation– and show that the DRL agent is successful at both tasks. At this point results are compared with the ones obtained with Particle Swarm Optimisation, certifying the advantage of DRL over more traditional gradient-free methods. We then consider a Transfer Learning scenario where the DRL agent is pre-trained using NeuralFoil [33], a faster-yet-less-accurate neural network surrogate model of XFoil built using domain knowledge. After an initial pre-training phase, we subsequently fine-tune the agent’s performance through transfer learning, incorporating direct training with XFoil. We compare four different transfer learning strategies and demonstrate how this combined approach can be successfully implemented, achieving accuracy levels comparable to those obtained through direct DRL training with XFoil, but at a reduced computational cost.
@@
The rest of the paper is organised as follows. In Section 2, we introduce the whole methodology, that combines a refined reward function for multi-objective optimisation, and using transfer learning coupled with reinforcement learning for airfoil shape optimisation. In Section 3, we report the results, comparing DRL agents with PSO (both for aerodynamic and aerodynamic/structural shape optimisation), trained with vs without transfer learning. Finally, in Section 4, we conclude.
2 Methodology
2.1 Problem under study in a nutshell
@@
We consider the two-dimensional (2D) optimisation of an airfoil shape under fixed operating conditions. The conditions chosen for the study are an Angle of Attack degrees, a Mach number , and a Reynolds number . These represent a typical aerodynamic scenario for subsonic flow, allowing us to investigate the optimisation of airfoil shapes under conditions that are common in practical applications. In general terms, we shall consider an airfoil shape to be ‘better’ than another one when the aerodynamic flow around the airflow shows an enhanced lift-to-drag ratio. Accordingly, the objective function to be maximised is such ratio, which is computed by applying an aerodynamic solver to any given airfoil geometry. In the next sections, we detail our proposed geometric optimisation framework and its specifications. We introduce a Reinforcement Learning paradigm which is enhanced by a Transfer Learning strategy where the agent suitably switches between using a fast, surrogate aerodynamic solver (NeuralFoil [33]) to a slightly slower yet more accurate one (XFoil [34]) as the method to assess the aerodynamic flow and to compute the agent’s reward. XFoil is a classic model –based on a panel method coupled with a viscous boundary layer formulation– designed for subsonic flows that calculates the pressure distribution on the airfoil and subsequently derives the lift and drag characteristics. It is faster than performing full DNS integration, but is still considered as an actual aerodynamic solver. NeuralFoil on the other hand –implemented in a Python-based library– is a ML surrogate model that can provide representation of the viscous, compressible airfoil aerodynamics for nearly any airfoil, with control surface deflections, across a 360-degree angle of attack, at any Reynolds number. NeuralFoil has been trained with millions of XFoil simulations, enabling rapid calculations of aerodynamic characteristics, albeit with a slight trade-off in accuracy. Unlike XFoil, Neuralfoil avoids non-convergence issues.
2.2 The basic DRL optimisation framework
@@
2.2.1 RL specs I: environment, state space and action definition
@@
coordinates of the points over the airfoil. However, this choice is delicate as it may lead the agent to explore unusual or even unphysical shapes. A better choice is to use the so-called Class-Shape Transformation (CST) parameters [35], a widely-used technique to parametrise geometries based on the curvature of the shape along the chord. This parameterization indeed generates more realistic shapes and reduces the dimensionality of the design space, allowing for a more efficient optimisation.
The implementation of such environment was done using the Gymnasium Python library [36], a maintained fork of the OpenAI Gym library, which provides a flexible framework for defining and managing RL environments.
State space. The state space spans a total of 18 scalar parameters , where each parameter is bounded in a concrete range . These 18 parameters are extracted from the CST representation, such that eight of them characterise the upper surface, eight of them the lower surface, one for the leading-edge weight, and one for the trailing-edge thickness. Different configuration values of these 18 parameters within their ranges correspond to different airfoil geometries.
@@
where EpisodeMaxLength is a positive integer defined in Sec.2.2.2. Accordingly, updating the state based on the action follows
@@
2.2.2 RL specs II: Episode definition
A single step of the DRL agent consists in, given an action, updating the airfoil geometry according to Eq. 2, assessing how this new geometry improves (or not) the aerodynamic efficiency and checking if the termination condition is met. We have decided that this condition will be reached when a maximum number of steps has been completed or if the aerodynamic solver does not converge.
@@ 2.2.2 RL specs II: Episode definition
Once an episode is defined, we need to define how the start of a new episode is carried out. A standard approach is to fix the initial state (to a certain airfoil geometry, e.g. naca0012), such that at the beginning of each episode the state of the environment is always reset to the same airfoil. This is, for example, the choice in [18]. Some other studies [19] in turn decide to reset to a different airfoil geometry, the rationale being that in this way the agent is capable of learning to optimise starting from a variety of shapes, i.e. the learned policy is more universally applicable. Here we chose the latter strategy, so that at the beginning of an episode the environment is chosen at random from a pre-compiled set of 20 naca airfoils listed in the Appendix A.1.
2.2.3 RL Specs III: Reward function
@@
Once the episode has been defined, we need to specify the reward function, which estimates the quality of any given action. As a starting point, we consider the lift-to-drag ratio , where both aerodynamic coefficients are computed from the aerodynamic solver. As previously mentioned, in this work we will use two solvers: XFoil and NeuralFoil.
@@
In addition to calculating lift and drag, Neuralfoil also returns a ‘confidence’ value: a scalar that indicates how confident the surrogate model is after making the lift and drag predictions. In order to penalise weird geometries or geometries whose computed aerodynamic properties have large uncertainty, we will aim to maximise , i.e. we aim at finding geometries that at the same time are improving the lift-to-drag ratio and keeping the confidence high. However, when using solely as reward the agent was not able to converge, that is why after looking [18] the reward was built with differences of the form
At this point, we need to include an extra-term in the reward –making the optimisation problem a multi-objective one– to model the effect of preserving the maximum thickness mt of the airfoil, the latter being a structural constraint of the airfoil. The rationale is to find optimal geometries from the point of view of their aerodynamic characteristics that, at the same time, keep constant (as much as possible) the maximum thickness, i.e. , so that the quotient remains close to 1. To penalise geometries that take such quotient away from one, we then define a Gaussian regularization
@@
where is an hyperparamenter that controls the strength of the nonlinear penalisation (see Fig. 2 for an illustration of how the values of penalises geometries with non-conserved maximum thickness, for different values of ). This Gaussian regularisation is inspired by computational fluid mechanics mesh optimisation methods that minimise error while maintaining low computational cost [37, 38]. Finally, we can define the reward at step as
Note that when we use XFoil then is not defined, in that case we set in Eq. 4.
Similarly, single-objective optimisation (where maximum thickness is not forced to be preserved) corresponds to the special case .
2.2.4 RL specs IV: the agent

### v2

# Transfer learning-enhanced deep reinforcement learning for aerodynamic airfoil optimisation subject to structural constraints

@@ Abstract
The main objective of this paper is to introduce a transfer learning-enhanced deep reinforcement learning (DRL) methodology that is able to optimise the geometry of any airfoil based on concomitant aerodynamic and structural integrity criteria. To showcase the method, we aim to maximise the lift-to-drag ratio while preserving the structural integrity of the airfoil –as modelled by its maximum thickness– and train the DRL agent using a list of different transfer learning (TL) strategies. The performance of the DRL agent is compared with Particle Swarm Optimisation (PSO), a traditional gradient-free optimisation method. Results indicate that DRL agents are able to perform purely aerodynamic and hybrid aerodynamic/structural shape optimisation, that the DRL approach outperforms PSO in terms of computational efficiency and aerodynamic improvement, and that the TL-enhanced DRL agent achieves performance comparable to the DRL one, while further saving substantial computational resources.
I Introduction
@@
Airfoil design is key to improving the performance and efficiency of various aerodynamic systems, such as aircraft wings, turbine blades, and wind turbines Lu (2022). Such design often predicates on (i) achieving an optimal balance between two aerodynamic forces –where one aims to maximise the aerodynamic lift and minimise the aerodynamic drag– while (ii) maintaining structural integrity Bak et al. (2014). The aerodynamic balance is often explicitly modelled as the lift-to-drag ratio , where and are the lift and drag coefficients. The optimisation problem can be formally stated as maximising an objective function , with representing the airfoil design parameters.
@@
The two common strategies for airfoil shape optimisation involve, respectively, using gradient-based or gradient-free methods. The former methods rely on the computation of to iteratively improve the design Skinner and Zare-Behtash (2018); et al. (2019). This approach is computationally efficient –particularly with high-dimensional design spaces–, making it a popular choice in aerodynamics as highlighted by comparative studies Secanell and Suleman (2005); Lyu and Martins (2014). The well-known adjoint method, introduced in Fluid Dynamics by the seminal work of Pironneau Pironneau (1974), is one prominent example of a gradient-based technique; requiring a computational effort that scales independently of the number of design variables. Subsequently, this method was refined by the so-called discrete adjoint method, becoming a widely used choice in this field Frank and Shubin (1992); Jameson, Martinelli, and Pierce (1998). While gradient-based methods are generally efficient and very popular in aerodynamic shape optimisation, they only have provable local convergence, and thus for non-convex objective functions their performance is sensitive to the initial design point Jichao, Xiaosong, and R.R.A.Martins (2022) and function continuity Skinner and Zare-Behtash (2018).
@@
On the other hand, gradient-free methods offer an alternative to complex, nonlinear optimisation problems that often appear in aerodynamic design Skinner and Zare-Behtash (2018); Viquerat et al. (2021). Despite their higher computational cost compared to gradient-based methods, they are more robust at finding global optima and handling challenging objective functions Skinner and Zare-Behtash (2018); et al. (2019). Popular gradient-free methods deployed in the context of airfoil shape optimisation
include Genetic Algorithms (GA) Bizzarrini et al. (2011) or
@@
Particle Swarm Optimisation (PSO) Wang, Zhang, and Chen (2011); Fourie and Groenwold (2002); Nejat, Mirzabeygi, and Panahi (2014). As advanced, these methods are often limited by their computational demands and can be inefficient with very large design spaces Lyu and Martins (2014). To partially bridge these issues, some recent efforts to improve the computational efficiency of gradient-free methods include, e.g., parallel implementations, which can make GAs viable even for intensive applications such as those in aerospace engineering Wang, Periaux, and Sefrioui (2002).
@@
Recently, deep reinforcement learning (DRL) Arulkumaran et al. (2017) has emerged as a promising (gradient-free) tool for shape optimisation Dussauge et al. (2023); Lou et al. (2023). DRL methods have demonstrated remarkable capabilities in various domains, particularly when dealing with complex high-dimensional data et al. (2019). Their impact in fluid dynamics studies and aeronautical applications belongs to a recently but rapidly expanding research field that encompasses both general-purpose ML-based fluid mechanics Brunton, Noack, and Koumoutsakos (2020); Vinuesa, Brunton, and McKeon (2023); Vinuesa and Brunton (2022), and their application to several aeronautical problems Le Clainche et al. (2023); Brunton et al. (2021) with clear industrial impact Lacasa et al. (2025).
@@
In the task of airfoil optimisation, a significant drawback of training DRL algorithms lies in the computational cost associated with solving the aerodynamic flow problem around the airfoil for a concrete shape: while the most accurate approach to solve the aerodynamic flow involves computational fluid dynamics (CFD) techniques, its intrinsic high computational cost requires opting for alternative, approximate solvers. For instance, some recent studies employ panel methods such as XFoil Dussauge et al. (2023); Lou et al. (2023) to approximate the aerodynamic performance of airfoils, while some other studies Liu et al. (2024, 2023) focus on further reducing optimisation costs by constructing surrogate models instead. Although the latter offers faster training times and smoother reward variations with respect to the agent’s actions –facilitating the learning process–, these approximate models come at the price of a decrease in accuracy.
@@
Interestingly, one significant advantage of recent deep learning methods in optimisation tasks is their potential to enhance their learning efficiency through so-called Transfer Learning Yan et al. (2019); Noda et al. (2023): the possibility of pre-training a given model on one task and subsequently fine-tuning it for a related but different task. Transfer learning could also be developed in the context of DRL, accelerating the agent’s learning process by reusing elements such as policies or feature representations and effectively reducing the need for extensive exploration, a particularly valuable trait in aerodynamic optimisation.
To the best of our knowledge this approach has been seldom explored, with a few encouraging exceptions for missile control surface and airfoils Yan et al. (2019); Bhola et al. (2023). The novelty we bring with respect to these two works is a greater reduction of the cost when transfer learning is applied, and the thorough implementation of a suite of transfer learning-enhanced DRL framework for aerodynamic optimisation subject to structural constraints.
@@
Additionally, note that most studies on airfoil shape optimisation focus solely on maximising aerodynamic efficiency Liu et al. (2024); Dussauge et al. (2023); Lou et al. (2023); Yan et al. (2019); Liu et al. (2025), without explicitly considering important structural integrity constraints Buckley, Zhou, and Zingg (2010) related, e.g., to the airfoil’s area or thickness. Accordingly, there is a need to develop shape optimisation frameworks that aim to maximise the aerodynamic efficiency while preserving the structural integrity of the airfoil. This, we argue, can be interpreted as an example of multi-objective optimisation Liu et al. (2024), which in practice can be reduced to a single-objective one by suitable scalarisation.
@@
To bridge these gaps, here we aim to explore the performance of DRL (enhanced with different possible Transfer Learning strategies) as a gradient-free framework for airfoil shape optimisation that considers the concomitant role of aerodynamic efficiency and structural integrity. To that aim, we define a flexible reward function that combines (i) the lift-to-drag ratio of the flow around the airfoil and (ii) the airfoil’s maximum thickness. We initially train our DRL agent using XFoil –both for aerodynamic optimisation alone, and combined aerodynamic shape optimisation with structural integrity preservation– and show that the DRL agent is successful at both tasks. At this point results are compared with the ones obtained with Particle Swarm Optimisation, certifying the advantage of DRL over more traditional gradient-free methods. We then consider a Transfer Learning scenario where the DRL agent is pre-trained using NeuralFoil Sharpe and Hansman (2025), a faster-yet-less-accurate neural network surrogate model of XFoil built using domain knowledge. After an initial pre-training phase, we subsequently fine-tune the agent’s performance through transfer learning, incorporating direct training with XFoil. We compare four different transfer learning strategies and demonstrate how this combined approach can be successfully implemented, achieving accuracy levels comparable to those obtained through direct DRL training with XFoil, but at a reduced computational cost.
@@
The rest of the paper is organised as follows. In Section II, we introduce the whole methodology. We start by formalizing the original multi-objective optimisation problem where aerodynamic and structural integrity criteria yield two different objective functions, and we suitably combine these into a reward function which effectively scalarises the problem into the framework of single-objective optimisation. Then, we introduce the reinforcement learning method and enhance it with several transfer learning strategies. In Section III, we report the results, comparing DRL agents with PSO (both) for pure aerodynamic shape optimisation and aerodynamic/structural integrity optimisation, trained with vs without transfer learning. Finally, in Section IV, we conclude.
II Methodology
II.1 Problem under study in a nutshell
We consider the two-dimensional geometric optimisation of an airfoil shape under fixed operating conditions. The conditions chosen for the study are an Angle of Attack degrees, a Mach number , and a Reynolds number . These represent a typical aerodynamic scenario for subsonic flow, allowing us to investigate the optimisation of airfoil shapes under conditions that are common in practical applications. In general terms, we shall consider an airfoil shape to be ‘better’ than another one when the aerodynamic flow around the airflow shows an enhanced lift-to-drag ratio. Accordingly, from an aerodynamic perspective the objective function to be maximised is such ratio, which is computed by applying an aerodynamic solver to any given airfoil geometry. Note, however, that aiming to optimise this ratio alone can yield geometries whose structural integrity can be compromised. Accordingly, in what follows we will also include structural constraints when we build the reward function.
In the next subsections, we detail our proposed geometric optimisation framework and its specifications. We will introduce a Reinforcement Learning paradigm which we enhance by incorporating a Transfer Learning strategy, where the agent suitably switches between using a fast, surrogate aerodynamic solver (NeuralFoil Sharpe and Hansman (2025)) to a slightly slower, yet more accurate one (XFoil Drela (1989)) as the method to assess the aerodynamic flow and to compute the agent’s reward. XFoil is
@@ II.1 Problem under study in a nutshell
a classic model –based on a panel method coupled with a viscous boundary layer formulation– designed for subsonic flows that calculates the pressure distribution on the airfoil and subsequently derives the lift and drag characteristics. It is faster than performing full DNS integration, but is still considered as an actual aerodynamic solver. NeuralFoil on the other hand –implemented in a Python-based library– is a ML surrogate model that can provide representation of the viscous, compressible airfoil aerodynamics for nearly any airfoil, with control surface deflections, across a 360-degree angle of attack, at any Reynolds number. NeuralFoil has been trained with millions of XFoil simulations, enabling rapid calculations of aerodynamic characteristics, albeit with a slight trade-off in accuracy. Unlike XFoil, NeuralFoil avoids non-convergence issues. While training the agent with NeuralFoil yields poorer results than when the agent is trained with Xfoil (see Table 5), we will show that we can successfully leverage NeuralFoil to reach optimal results in the context of Transfer Learning enhancement.
As a final note, the developed DRL geometric optimisation framework is available through the open source library pyLOM Eiximeno et al. (2025).
II.2 The basic DRL optimisation framework
@@
II.2.1 RL specs I: environment, state space and action definition
@@
coordinates of the points over the airfoil. However, this choice is delicate as it may lead the agent to explore unusual or even unphysical shapes. A better choice is to use the so-called Class-Shape Transformation (CST) parameters Kulfan (2008), a widely-used technique to parametrise geometries based on the curvature of the shape along the chord. This parameterisation indeed generates more realistic shapes and reduces the dimensionality of the design space, allowing for a more efficient optimisation.
The implementation of such environment was done using the Gymnasium Python library Towers et al. (2024), a maintained fork of the OpenAI Gym library, which provides a flexible framework for defining and managing RL environments.
State space. The state space spans a total of 18 scalar parameters , where each parameter is bounded in a concrete range . LowerBound and UpperBound are two vectors that define the boundaries of the state space and are defined in Appendix A. These 18 parameters are converted to an airfoil using CST representation, such that eight of them characterise the upper surface, eight of them the lower surface, one for the leading-edge weight, and one for the trailing-edge thickness. Different configuration values of these 18 parameters within their ranges correspond to different airfoil geometries.
@@
where EpisodeMaxLength is a positive integer that represents the maximum permitted length of an episode, defined with more detail in Sec. II.2.2. This allows the agent to potentially explore both limits of the state space in a single episode, regardless of the initial state. Accordingly, updating the state based on the action follows
@@
II.2.2 RL specs II: Episode definition
A single step of the DRL agent consists in, given an action, updating the airfoil geometry according to Eq. 2, assessing how this new geometry improves (or not) the aerodynamic efficiency and checking if the termination condition is met. We have decided that this condition will be reached when a maximum number of steps has been completed or if the aerodynamic solver, in this case XFoil, does not converge.
@@ II.2.2 RL specs II: Episode definition
Once an episode is defined, we need to define how the start of a new episode is carried out. A standard approach is to fix the initial state (to a certain airfoil geometry, e.g. naca0012), such that at the beginning of each episode the state of the environment is always reset to the same airfoil. This is, for example, the choice in Dussauge et al. (2023). Some other studies Lou et al. (2023) in turn decide to reset to a different airfoil geometry, the rationale being that in this way the agent is capable of learning to optimise starting from a variety of shapes, i.e. the learned policy is more universally applicable. Here we chose the latter strategy, so that at the beginning of an episode the environment is chosen at random from a pre-compiled set of 20 naca airfoils listed in the Appendix A.
II.2.3 RL Specs III: Reward function
@@
Once the episode has been defined, we need to specify the reward function, which estimates the quality of any given action. For pedagogical reasons, we now outline the construction of such reward function in a step-by-step way. As a starting point, we consider the maximisation of the lift-to-drag ratio –where both aerodynamic coefficients are computed from the aerodynamic solver–, which yields a purely aerodynamic shape optimisation.
@@
As previously mentioned, in this work we will use two solvers: XFoil and NeuralFoil. Observe that, in addition to calculating lift and drag coefficients, NeuralFoil also returns a ‘confidence’ value: a scalar that indicates how confident the surrogate model is after making the lift and drag predictions. In order to penalise weird geometries or geometries whose computed aerodynamic properties have large uncertainty, we will aim to maximise , i.e. we aim at finding geometries that at the same time are (i) improving the lift-to-drag ratio and (ii) keeping the confidence high. With this term , we effectively mitigate the risk of the agent relying on potentially inaccurate predictions in regions where NeuralFoil is known to perform poorly.
Now, since in RL the agent tries to maximise the cumulative reward Dussauge et al. (2023), the actual reward function needs to take a differential form, e.g.
So far, as previously mentioned the reward function denotes a purely aerodynamic optimisation. So at this point we consider as a second objective function the needs to preserve the structural integrity of the airfoil. As a simply proxy, we consider the airfoil’s maximum thickness mt, and our goal is to preserve as much as possible this structural property, i.e. we aim to minimise the mismatch between the initial and final thicknesses. We formalise this by using a Gaussian kernel , such that at step , the kernel reads
@@
where is the ratio between the airfoil’s maximum thickness at step and the initial one, and is an hyperparameter that controls the strength of the nonlinear penalisation. This Gaussian regularisation is inspired by computational fluid mechanics mesh optimisation methods that minimise error while maintaining low computational cost Huergo, Rubio, and Ferrer (2024); Huergo et al. (2024). For a fixed , simply penalizes those airfoils whose maximum thickness is not preserved (see Fig. 2 for an illustration of how the values of penalises geometries with non-conserved maximum thickness, for different values of ). As such we now have to maximize the term given in Eq. 3 and, independently, maximize in Eq. 4. This, in essence, is a multi-objective optimisation problem: we aim to find airfoil geometries that maximise the aerodynamic properties and maximally preserve the initial maximum thickness (a proxy for preserving the structural integrity). Now, instead of treating the aerodynamic and the structural aspects independently, we choose to scalarize the multi-objective problem by combining the aerodynamic and the structural terms into a single reward function. Finally, we can define such reward at step as
Note that when we use XFoil then is not defined, in that case we set in Eq. 5. Since we now have a single reward function, observe that from a mathematical point of view we have approximated a multi-objective optimisation problem with a simpler, single-objective one which balances aerodynamic and structural integrity constraints. The Gaussian kernel effectively acts as a structural regularisation term.
In what follows, we will distinguish the special case (yielding which will be called the purely aerodynamic optimisation, from the more general case where structural integrity enters in the reward function. This latter scenario will be called the hybrid aerodynamic/structural optimisation case, although we need to make clear at this point that optimisation in the structural dimension is simply aimed at minimising the change in maximum thickness.
II.2.4 RL specs IV: the agent